Suppose you own an outdoor recreation company and you want to purchase all-terrain vehicles (ATVs) for your summer business and snowmobiles for your winter business. Your budget for new vehicles this year is $240,000. ATVs cost $8000 each and snowmobiles cost $12,000 each.
a. Draw a budget line for your purchase of new vehicles.
b. What is the opportunity cost of one ATV?
c. What is the opportunity cost of one snowmobile?
1
Expert's answer
2020-05-07T10:29:06-0400


a. Budget line is represented on the graph:

 


b. Opportunity cost of one ATV is 8 000/12 000 = 2/3 snowmobiles, as money we spend to buy 1 ATV can be alternatively used to buy 2/3 of snowmobile.

 

c. Analogously, opportunity cost of one snowmobile is 12 000/ 8 000=1.5 ATW